Solve the initial value problem.


y′′+y=1, y(0)=2, y′(0)=7.


1
Expert's answer
2022-03-29T11:47:45-0400

"y''+\\left( y-1 \\right) =0\\\\\\left( y-1 \\right) ''+\\left( y-1 \\right) =0\\\\\\lambda ^2+1=0\\Rightarrow \\lambda =\\pm i\\\\y-1=C_1\\cos t+C_2\\sin t\\\\y=1+C_1\\cos t+C_2\\sin t\\\\\\left\\{ \\begin{array}{c}\ty\\left( 0 \\right) =2\\\\\ty'\\left( 0 \\right) =7\\\\\\end{array} \\right. \\Rightarrow \\left\\{ \\begin{array}{c}\t1+C_1=2\\\\\tC_2=7\\\\\\end{array} \\right. \\Rightarrow \\left\\{ \\begin{array}{c}\tC_1=1\\\\\tC_2=7\\\\\\end{array} \\right. \\\\y=1+7\\cos t+\\sin t"